# Telco-oRAG: Optimizing Retrieval-augmented Generation for Telecom Queries via Hybrid Retrieval and Neural Routing

### v1

@@ I Introduction
However, vanilla LLMs rely solely on their internal representations and learned parameters to generate text, and they struggle in specialized domains such as telecommunications, where queries often involve domain-specific terminology, implicit standards knowledge, and subtle contextual dependencies. Even advanced models like GPT-4 exhibit limited reliability when tasked with answering questions related to 3GPP specifications [1]. Addressing these limitations is essential to fully realize the promise of AI-native networks.
@@ I-A Related Works
Fine-tuning offers an effective means to inject domain expertise into LLMs [2]. However, it requires substantial computational resources and is prone to overfitting when trained on limited data [3]. Moreover, fine-tuning exhibits reduced flexibility in rapidly evolving domains, which leads LLM vendors to periodically retrain their models to keep up with updates in specialized domains [4]. This makes any LLM obsolete a few months after it is released and results in unstable performance from one release to another [5], making LLM benchmarking very challenging.
Recent research has explored parameter-efficient strategies to mitigate the high computational costs of fine-tuning LLMs for domain-specific tasks [6]. Conventional fine-tuning requires storing large gradient states and domain-specific representations, leading to significant memory requirements. Techniques such as adapter layers [7], prefix-tuning [8], low-rank adaptation (LoRA) [9], or combining low-rank updates with quantization of the frozen base model (QLoRA) [10], have been proposed to minimize the number of trainable parameters while preserving model performance. Importantly, a late work has shown that LoRA does not inject sufficient additional knowledge to adapt the LLMs to the telecommunications field [11].
Recent efforts in telecom-specific fine-tuning include TelecomGPT [12], which leverages continual pre-training, instruct tuning with supervised fine-tuning (SFT), and alignment tuning with direct Preference Optimization (DPO) [13]. Each of these phases uses a specifically designed telecom dataset constructed using on public documents. Moreover, Tele-LLMs is a series of LLMs [11], ranging from 1B to 8B parameters developed using continual pre-training on a dedicated telecom dataset (Tele-Data).
@@
RAG has emerged as an alternative pathway by decoupling the injection of knowledge from the training of model parameters. By retrieving the relevant context from external knowledge sources at inference time, RAG reduces reliance on static model parameters and allows for more efficient adaptation to new information [14]. This approach has proven particularly effective in knowledge-intensive tasks, improving factual accuracy while avoiding the computational costs of periodic fine-tuning [15].
In recent years, the research community has proposed several advancements to improve RAG, which can be classified as follows [16]: 1) input enhancement, including query transformation and data augmentation; 2) retriever enhancement, such as chunk optimization, retriever finetuning, hybrid retrieval, and re-ranking; 3) generator enhancement, which includes prompt engineering, decoding tuning, and generator finetuning; 4) result enhancement, which allows RAG output rewrite; 5) RAG pipeline enhancement, including adaptive retrieval and iterative RAG.
The integration of RAG into telecommunications systems has shown substantial value [17]. Indeed, telecommunications systems support tools such as chatbots that streamline the access of professionals to standards, accelerating research, development and improving regulatory compliance. Few RAG frameworks have been developed to address the complexities of technical standards and their rapid evolution. For instance, TelecomRAG [18] utilizes a knowledge base built from 3GPP Release 16 and Release 18 specification documents to provide responses to telecom standard query. Importantly, in TelecomRAG, the retriever transform each new query using the history of past queries and responses. Telco-RAG [19] is an open-source framework designed to handle the specific needs of queries about telecommunications standards, by integrating domain-specific input, retrieval, and generator enhancements. Additionally, Chat3GPP [20] offers another open-source RAG tailored for 3GPP specifications, combining chunking strategies, and re-ranking. More recently, researchers have proposed CommGPT [21], a telecom specialized LLM constructed using dedicated continuing pre-training dataset and instruction fine-tuning dataset. The fine-tuned model is then combined with Knowledge Graph (KG) [22] and RAG to assist CommGPT in generating more precise and comprehensive responses. Although these approaches introduce significant novelties, leading to notable results, they all lack a means to simultaneously provide LLMs 1) a macroscopical vision of the topics related to the user query, 2) up-to-date information, and 3) precise understanding of domain-specific technical terms and abbreviations.
@@ I-B Main Contributions
    We design an neural Network (NN) router that selects the most appropriate sources of information for answering to user queries. This approach makes Telco-oRAG scalable with respect to future knowledge bases describing new technologies and products. Our results show that the proposed NN router is more accurate than off-the-shelf classifiers and that it reduces memory usage of baseline RAG by 45%.
@@
This journal significantly extends our earlier conference manuscript [19]: we introduce (i) an improved query refinement, (ii) a web retrieval module, (iii) a methodological optimization of the RAG hyperparameters, and (iv) an extensive set of numerical evaluations.
@@ II Preliminaries on RAG
The embedding model—e.g., OpenAI’s text-embedding-3-large [23]—encodes the semantic content of each chunk into a continuous vector space, enabling efficient similarity-based retrieval.
@@ III-A User Query Refinement
[…]
@@ III-A2 Glossary-enhancement
To achieve this goal, we construct two domain dictionaries based on the “Vocabulary for 3GPP Specifications” [24]:
@@ III-C2 Reducing the RAM requirements of Telco-oRAG
To address this issue, we recall that the 3GPP standards categorize specifications into 18 distinct series numbered from 21 to 38 [25]. Each series provides the technical details of a specific aspect of 3GPP standards (e.g., radio access, core network components, security). To filter out irrelevant information and reduce the RAG requirements on random-access memory (RAM) resources, we have developed an NN router tailored to infer the 3GPP series that contains required information for providing the correct answer to the user queries.
@@
The architecture of the proposed NN router is illustrated in Fig. 3. The process begins by constructing a high-level summary for each of the 18 3GPP specification series [25], using a dedicated LLM. These summaries are short textual descriptions that capture the thematic content of each series (e.g., radio protocols, access, core network). An example of one of the summaries, the one related to the Requirements series, is depicted in Box 2. Each summary is then embedded using the text-embedding-3-large model, resulting in a 1024-dimensional representation per series.
@@
To train the NN router, we constructed a dedicated dataset of 30,000 questions extracted from 500 documents in 3GPP Release 18, with each question labeled by its originating series. This approach avoids overfitting and ensures a robust evaluation of the Telco-oRAG pipeline [26].
@@ III-D Prompt Engineering
Prompt formulation plays a critical role in RAG, as it determines how effectively the language model incorporates retrieved context to generate a correct and concise answer [27]. In this work, we design a structured prompt format optimized for clarity and alignment with the user query.
The final prompt in Telco-oRAG follows a dialogue-oriented structure, which has been shown to improve LLM performance in multi-hop and context-heavy reasoning tasks [28]. The prompt begins with the refined query $\mathcal{Q}^{+}$, followed by the validated context retrieved from both 3GPP specifications and web documents.
@@ IV Experimental Results
In this section, we evaluate the performance of Telco-oRAG in answering questions related to telecommunications standards. To assess the effectiveness of our approach, we compare Telco-oRAG against several benchmarks, including current state-of-the-art LLMs and Telco-RAG [19]. In the following, if not indicated differently, Telco-oRAG uses GPT-3.5 as base model in each of its LLM components and its NN router filters k=5 3GPP series to extract domain-specific content from 3GPP Release 18 documents. The accuracy reported in each experiment is measured as the percentage of correct answers.
@@
  • Optimization dataset: A set of 2000 MCQs generated using the methodology proposed in [1], based on 3GPP Release 18 documents.
@@
  • 3GPP Standard dataset: A curated set of 1840 MCQs extracted from the TeleQnA [1], targeting 3GPP standard queries.
  • TeleQnA: A dataset including 10000 telecom-related MCQs [1].
@@ IV-A3 Embedding Model Selection
The embedding model converts queries and documents into vector representations. We compare text-embedding-ada-002 and text-embedding-3-large, which differ in embedding dimensionality and semantic encoding capacity  [23]. The latter model incorporates Matryoshka Representation Learning [29], which improves performance while compressing vector representations. With a fixed embedding dimension of 1024, text-embedding-3-large consistently outperforms text-embedding-ada-002, achieving an average accuracy improvement of 2.29% on the optimization dataset. Consequently, we adopt text-embedding-3-large as the default embedding model for Telco-oRAG.
@@ IV-A4 Indexing Strategy Selection
We consider three FAISS-based [30] similarity search methods: IndexFlatIP, IndexFlatL2, and IndexHNSW. Specifically, IndexFlatIP uses inner product, IndexFlatL2 computes exact Euclidean distance, and IndexHNSW provides approximate nearest-neighbor search with sublinear query time. As mentioned in Section II, cosine similarity is equivalent to the inner product when using $\ell_{2}$-normalized embeddings, making both IndexFlatIP and IndexFlatL2 theoretically suitable for retrieval. Our empirical results have confirmed this equivalence in practice, with both strategies yielding nearly identical rankings. However, despite only marginal differences in accuracy, IndexFlatIP has outperformed IndexFlatL2 in 80% of our experiments. By contrast, IndexHNSW, an approximate method optimized for speed, leads to substantial accuracy degradation due to its non-exact nature.
@@ IV-D Is Online All You Need?
In this section, we compare web search and classic RAG with respect to different types of telecom questions and analyze the gain brought by combining them together, i.e., in Telco-oRAG.
In Figure 7, using TeleQnA, we compare four LLM configurations based on GPT-3.5:
@@
These results highlight that the Baseline model achieves limited accuracy, in particular for standard overview and standard specification questions. However, both the online retrieval and Telco-RAG significantly enhance the baseline model performance in these domains as well as for the lexicon MCQs. Notably, the online retrieval outperforms Telco-RAG on standard overview while Telco-RAG achieves better accuracy than the online retrieval on standard specifications. Finally, Figure 7 shows that Telco-oRAG—combining domain-specific and online retrieval—achieves the highest accuracy across all categories.
@@
On 3GPP Standard Overview MCQs, GPT-3.5 and GPT-4 achieve 55.7% and 68.2% of accuracy, respectively. Augmenting GPT-3.5 with web search increases accuracy its 70.3%. Telco-RAG again outperforms the other baseline model, achieving 78.9% of accuracy. Finally, Telco-oRAG provides the best performance with 83.9% of accuracy, leading to more than 25% and 15% improvement over GPT-3.5 and GPT-4.
It is important to note that the performances in Figure 8 are not the same as the one for the standard specifications and standard overview in Figure 8, as 1) TeleQnA includes MCQs from different standards while the 3GPP evaluation dataset has questions only related to 3GPP documents, 2) in our experiments, Telco-RAG and Telco-oRAG leverage a database composed exclusively by 3GPP Rel. 18 documents, which makes them particularly effective on the 3GPP evaluation dataset.

### v2

@@ I Introduction
However, vanilla LLMs rely solely on their internal representations and learned parameters to generate text, and they struggle in specialized domains such as telecommunications, where queries often involve domain-specific terminology, implicit standards knowledge, and subtle contextual dependencies. Even advanced models like GPT-4 exhibit limited reliability when tasked with answering questions related to 3GPP specifications [maatouk2023teleqna]. Addressing these limitations is essential to fully realize the promise of AI-native networks.
@@ I-A Related Works
Fine-tuning offers an effective means to inject domain expertise into LLMs [gururangan2020don]. However, it requires substantial computational resources and is prone to overfitting when trained on limited data [thompson2020computational]. Moreover, fine-tuning exhibits reduced flexibility in rapidly evolving domains, which leads LLM vendors to periodically retrain their models to keep up with updates in specialized domains [IBM]. This makes any LLM obsolete a few months after it is released and results in unstable performance from one release to another [chen2023chatgptsbehaviorchangingtime], making LLM benchmarking very challenging.
Recent research has explored parameter-efficient strategies to mitigate the high computational costs of fine-tuning LLMs for domain-specific tasks [zhang2024instructiontuninglargelanguage]. Conventional fine-tuning requires storing large gradient states and domain-specific representations, leading to significant memory requirements. Techniques such as adapter layers [Mahabadi2021], prefix-tuning [li-liang-2021-prefix], low-rank adaptation (LoRA) [hu2021lora], or combining low-rank updates with quantization of the frozen base model (QLoRA) [Dettmers2023], have been proposed to minimize the number of trainable parameters while preserving model performance. Importantly, a late work has shown that LoRA does not inject sufficient additional knowledge to adapt the LLMs to the telecommunications field [maatouk2024telellmsseriesspecializedlarge].
Recent efforts in telecom-specific fine-tuning include TelecomGPT [zou2024telecomgpt], which leverages continual pre-training, instruct tuning with supervised fine-tuning (SFT), and alignment tuning with direct Preference Optimization (DPO) [rafailov2024directpreferenceoptimizationlanguage]. Each of these phases uses a specifically designed telecom dataset constructed using on public documents. Moreover, Tele-LLMs is a series of LLMs [maatouk2024telellmsseriesspecializedlarge], ranging from 1B to 8B parameters developed using continual pre-training on a dedicated telecom dataset (Tele-Data).
@@
RAG has emerged as an alternative pathway by decoupling the injection of knowledge from the training of model parameters. By retrieving the relevant context from external knowledge sources at inference time, RAG reduces reliance on static model parameters and allows for more efficient adaptation to new information [Lewis2020NIPS]. This approach has proven particularly effective in knowledge-intensive tasks, improving factual accuracy while avoiding the computational costs of periodic fine-tuning [izacard2022few].
In recent years, the research community has proposed several advancements to improve RAG, which can be classified as follows [zhao2024retrievalaugmentedgenerationaigeneratedcontent]: 1) input enhancement, including query transformation and data augmentation; 2) retriever enhancement, such as chunk optimization, retriever finetuning, hybrid retrieval, and re-ranking; 3) generator enhancement, which includes prompt engineering, decoding tuning, and generator finetuning; 4) result enhancement, which allows RAG output rewrite; 5) RAG pipeline enhancement, including adaptive retrieval and iterative RAG.
The integration of RAG into telecommunications systems has shown substantial value [piovesan2024telecom]. Indeed, telecommunications systems support tools such as chatbots that streamline the access of professionals to standards, accelerating research, development and improving regulatory compliance. Few RAG frameworks have been developed to address the complexities of technical standards and their rapid evolution. For instance, TelecomRAG [Yilma2025] utilizes a knowledge base built from 3GPP Release 16 and Release 18 specification documents to provide responses to telecom standard query. Importantly, in TelecomRAG, the retriever transform each new query using the history of past queries and responses. Telco-RAG [bornea2024telcorag] is an open-source framework designed to handle the specific needs of queries about telecommunications standards, by integrating domain-specific input, retrieval, and generator enhancements. Additionally, Chat3GPP [huang2025chat3gpp] offers another open-source RAG tailored for 3GPP specifications, combining chunking strategies, and re-ranking. More recently, researchers have proposed CommGPT [jiang2025], a telecom specialized LLM constructed using dedicated continuing pre-training dataset and instruction fine-tuning dataset. The fine-tuned model is then combined with Knowledge Graph (KG) [abusalih2021domainspecificknowledgegraphssurvey] and RAG to assist CommGPT in generating more precise and comprehensive responses. Although these approaches introduce significant novelties, leading to notable results, they all lack a means to simultaneously provide LLMs 1) a macroscopical vision of the topics related to the user query, 2) up-to-date information, and 3) precise understanding of domain-specific technical terms and abbreviations.
@@ I-B Main Contributions
    We design an neural Network (NN) router that selects the most appropriate sources of information for answering user queries. This approach makes Telco-oRAG scalable with respect to future knowledge bases describing new technologies and products. Our results show that the proposed NN router is more accurate than off-the-shelf classifiers and that it reduces memory usage of baseline RAG by 45%. Moreover, we analyse the failure cases of the proposed NN router and their impact on the overall performance.
@@
    We provide an end-to-end latency analysis of Telco-oRAG, breaking down per-stage costs and discussing latency-performance tradeoff of each component in Telco-oRAG.
  6. 6.
    We extend our evaluation to open-ended QnA using the LLM-as-judge method. These results confirm the value of Telco-oRAG that leads up to 42.8 percentage-points gains with respect to baseline models.
  7. 7.
@@
This journal significantly extends our earlier conference manuscript [bornea2024telcorag]: we introduce (i) an improved query refinement, (ii) a web retrieval module, (iii) a methodological optimization of the RAG hyperparameters, and (iv) an extensive set of numerical evaluations including latency analysis and performance evaluation on open-ended questions.
@@ II Preliminaries on RAG
The embedding model—e.g., OpenAI’s text-embedding-3-large [openai2023embeddings]—encodes the semantic content of each chunk into a continuous vector space, enabling efficient similarity-based retrieval.
@@
The folowing discussion mainly focuses on answering to questions related to 3GPP questions. This choice was done considering the importance of 3GPP in the telecommuni- cation domain. However, in our results, we show that even a Telco-oRAG implementation targeting 3GPP standards is able to answer to questions from other sources; moreover, extending this pipeline on other standards or documents does not need any architecture change and has as its main focus 3GPP Standards, both as questions and documents.
This was only a matter of choice, driven by the underperfomance of LLMs on this set of questions. Future research can test the full Telco-oRAG pipeling on other standards or documents without any arhitectural modifications. It only requires: (i) updating the embeddings database, (ii) updating abbreviations and technical terms, and (iii) updating the neural router training set (for instance considering the different ITU recommendations series) and its model accordingly.
@@ III-A User Query Refinement
$s=\text{``What role does PCRF play in QoS control?"},$
@@ III-A2 Glossary-enhancement
To achieve this goal, we construct two domain dictionaries based on the “Vocabulary for 3GPP Specifications” [3GPPTR21.905]:
@@ III-C2 Reducing the RAM requirements of Telco-oRAG
To address this issue, we recall that the 3GPP standards categorize specifications into 18 distinct series numbered from 21 to 38 [specsbyseries]. Each series provides the technical details of a specific aspect of 3GPP standards (e.g., radio access, core network components, security). To filter out irrelevant information and reduce the RAG requirements on random-access memory (RAM) resources, we have developed an NN router tailored to infer the 3GPP series that contains required information for providing the correct answer to the user queries.
@@
The architecture of the proposed NN router is illustrated in Fig. 3. The process begins by constructing a high-level summary for each of the 18 3GPP specification series [specsbyseries], using a dedicated LLM. These summaries are short textual descriptions that capture the thematic content of each series (e.g., radio protocols, access, core network). An example of one of the summaries, the one related to the Requirements series, is depicted in Box 2. Each summary is then embedded using the text-embedding-3-large model, resulting in a 1024-dimensional representation per series.
@@
To train the NN router, we constructed a dedicated dataset of 30,000 questions extracted from 500 documents in 3GPP Release 18, with each question labeled by its originating series. This approach avoids overfitting and ensures a robust evaluation of the Telco-oRAG pipeline [Gilardi2023].
@@ III-D Prompt Engineering
Prompt formulation plays a critical role in RAG, as it determines how effectively the language model incorporates retrieved context to generate a correct and concise answer [CHEN2025101260]. In this work, we design a structured prompt format optimized for clarity and alignment with the user query.
The final prompt in Telco-oRAG follows a dialogue-oriented structure, which has been shown to improve LLM performance in multi-hop and context-heavy reasoning tasks [Gao2019NeuralAT]. The prompt begins with the refined query $\mathcal{Q}^{+}$, followed by the validated context retrieved from both 3GPP specifications and web documents.
@@ IV Experimental Results
In this section, we evaluate the performance of Telco-oRAG in answering questions related to telecommunications standards. To assess the effectiveness of our approach, we compare Telco-oRAG against several benchmarks, including current state-of-the-art LLMs and Telco-RAG [bornea2024telcorag]. In the following, if not indicated differently, Telco-oRAG uses GPT-3.5 as base model in each of its LLM components and its NN router filters k=5 3GPP series to extract domain-specific content from 3GPP Release 18 documents. The accuracy reported in each experiment is measured as the percentage of correct answers.
@@
  • Optimization dataset: A set of 2000 MCQs generated using the methodology proposed in [maatouk2023teleqna], based on 3GPP Release 18 documents.
@@
  • 3GPP Standard dataset: A curated set of 1840 MCQs extracted from the TeleQnA [maatouk2023teleqna], targeting 3GPP standard queries.
  • TeleQnA: A dataset including 10000 telecom-related MCQs, based on research manuscripts and documents produced by different standard organization [maatouk2023teleqna].
@@ IV-A3 Embedding Model Selection
The embedding model converts queries and documents into vector representations. We compare text-embedding-ada-002 and text-embedding-3-large, which differ in embedding dimensionality and semantic encoding capacity  [openai2023embeddings]. The latter model incorporates Matryoshka Representation Learning [kusupati2022matryoshka], which improves performance while compressing vector representations. With a fixed embedding dimension of 1024, text-embedding-3-large consistently outperforms text-embedding-ada-002, achieving an average accuracy improvement of 2.29% on the optimization dataset. Consequently, we adopt text-embedding-3-large as the default embedding model for Telco-oRAG.
@@ IV-A4 Indexing Strategy Selection
We consider three FAISS-based [johnson2017faiss] similarity search methods: IndexFlatIP, IndexFlatL2, and IndexHNSW. Specifically, IndexFlatIP uses inner product, IndexFlatL2 computes exact Euclidean distance, and IndexHNSW provides approximate nearest-neighbor search with sublinear query time. As mentioned in Section II, cosine similarity is equivalent to the inner product when using $\ell_{2}$-normalized embeddings, making both IndexFlatIP and IndexFlatL2 theoretically suitable for retrieval. Our empirical results have confirmed this equivalence in practice, with both strategies yielding nearly identical rankings. However, despite only marginal differences in accuracy, IndexFlatIP has outperformed IndexFlatL2 in 80% of our experiments. We hypothesize that this advantage arises from FAISS optimized handling of inner product computations, which may benefit from lower numerical error accumulation compared to explicit $\ell_{2}$-distance evaluation. This difference, while subtle, proved consistent across diverse subsets, suggesting that implementation-level details can have a measurable effect even when methods are theoretically equivalent. By contrast, IndexHNSW, an approximate method optimized for speed, leads to substantial accuracy degradation due to its non-exact nature. Nevertheless, we acknowledge that IndexHNSW may remain attractive in large-scale or latency-sensitive deployments, where its speed offers a practical trade-off between retrieval quality and efficiency.
@@ IV-C The Memory Footprint of Telco-oRAG
Figure 8: NN router prediction quality per 3GPP series.
Figure 7 the distribution of the misclassified questions across 3GPP series. These results show that the errors exhibit some patterns such as 24 series and 28 series being frequently confused with 38 series or 23 series with 29 series. These series operate in adjacent or semantically related domains within 3GPP, often sharing overlapping technical terminology and collaborative work items.
In addition, Figure 4 reports the per–working-group prediction accuracy at Top-1, Top-3, and Top-5 levels. The results demonstrate that while several working groups achieve high reliability at Top-1, others exhibit substantially lower performance. Notably, a 50% accuracy threshold across working groups is only attained at the Top-5 level, thereby motivating our choice to propagate a Top-5 shortlist downstream in the pipeline. Finally, despite these misclassifications, our experiments (using GPT-5-mini) show that Telco-oRAG performance does not drop with respect a baseline model, even when all the 5 retrieved documents are wrong, highlight that our pipeline remain robust, thanks to the web retrieval stage.
@@ IV-D Is Online All You Need?
In this section, we compare web search and classic RAG with respect to different types of telecom questions and analyze the gain brought by combining them together, i.e., in Telco-oRAG. In Figure 9, using TeleQnA, we compare four LLM configurations based on GPT-3.5:
@@
These results highlight that the baseline GPT-3.5 model achieves limited accuracy, in particular for standard overview and standard specification questions. However, both the web retrieval and Telco-RAG significantly enhance the baseline model performance in these domains as well as for the lexicon MCQs. Notably, the online retrieval outperforms Telco-RAG on standard overview while Telco-RAG achieves better accuracy than the online retrieval on standard specifications. These results are due to the limited capability of Telco-RAG to answer to questions related to non-3GPP standards, as its embedding database is based on 3GPP Rel.18 documents. Finally, Figure 9 shows that Telco-oRAG—combining domain-specific and online retrieval—achieves the highest accuracy across all TeleQnA categories, most notably on ”Standards Overview” where it brings $\approx$10 pp in comparison to Telco-RAG.
@@
On 3GPP Standard Overview MCQs, GPT-3.5 and GPT-4 achieve 55.7% and 68.2% of accuracy, respectively. Augmenting GPT-3.5 with web search increases accuracy its 70.3%. Telco-RAG again outperforms the other baseline model, achieving 78.9% of accuracy. Finally, Telco-oRAG provides the best performance with 83.9% of accuracy, leading to more than 25% and 15% improvement over GPT-3.5 and GPT-4. z It is important to note that the performances in Figure 10 are not the same as the one for the standard specifications and standard overview in Figure 10, as 1) TeleQnA includes MCQs from different standards while the 3GPP evaluation dataset has questions only related to 3GPP documents, 2) in our experiments, Telco-RAG and Telco-oRAG leverage a database composed exclusively by 3GPP Rel. 18 documents, which makes them particularly effective on the 3GPP evaluation dataset.